# A Domain-specific Language and Architecture for Detecting Process Activities from Sensor Streams in IoT

### v1

@@ 22 In SM_1 sensor o6_valve
In Listing LABEL:lst:syntax-example1 we present a Radiant example for parts of the blood donation process (cf. Section 2.2). The start of the Sanitize hands activity (line 2) is detected when the load cell sensor in the central hygiene station (CHS) changes its state from low to high (lines 4–5). This pattern requires a discretization of the sensor values emitted from the load cell (in gramms) to more abstract states. In our setup, we assume that the bottle with the hand sanitizer is placed on the load cell and it has an initial weight which will decrease with every usage. We denote this state with slowly decreasing weights as low. When the healthcare worker applies pressure to dispense liquid from the sanitizer, we can observe a significant brief peak in the load cell’s values ($>$ 1000 g) in a value range we denote as high. Similarly, we detect the end of this activity based on a change from high to low (lines 6–8). Note that for both conditions, we specify a time window constraint of 30 seconds as we assume that the healthcare worker might use the dispenser several times to perform one instance of the Sanitize hands activity.
@@ 4.5 Code Generation
The core requirement of the target language and execution platform should be its ability to process sensor events from one or more event streams and detect specific patterns in the current events as well as in subsequent events. Complex event processing (CEP) is the most suitable technology here as it supports exactly these features [59]. As a concrete CEP platform, we decided for Siddhi which is part of the WSO2 stream processing platform [64]. Siddhi is a light-weight Java-based service optimized for high-velocity and high-throughput event processing. It features its own event processing language called StreamingSQL [58], which is the main target (host) language of our implemented code generator to translate the Radiant artefacts to. StreamingSQL supports all concepts required to detect patterns in sensor event streams w.r.t. the corresponding activity executions. However, as pointed out in [47], CEP languages are still too complex and require technical expertise to effectively write event processing applications. Here, following a model-driven engineering approach [47], Radiant introduces necessary abstractions to support domain experts with writing applications for the specific use case of activity detection. The Siddhi CEP platform includes a runtime system where the generated Siddhi apps can be directly deployed to and executed via a REST API, thus allowing seamless development workflows from composition with Radiant to execution and activity detection at runtime from live IoT sensor data. If necessary, domain experts can also inspect and modify the generated Siddhi apps before deployment in a web-based editor integrated with the Siddhi CEP platform.
@@
Listing LABEL:lst:siddhi-fragments presents exemplary fragments of a Siddhi app generated from the Radiant example for the Sort product activity in the production process (cf. Listing LABEL:lst:syntax-example2). We generate one app per activity that is defined in a Radiant application with the process name providing the execution context of an activity. Note that this leads to two separate Siddhi apps being created for Listing LABEL:lst:syntax-example2. As discussed before, their ordering within one application does not play a role. For each station with its sensors used in a Radiant application, we additionally retrieve the IoT configuration parameters from the YAML file (cf. Section 4.3) during code generation to add these configurations to the beginning of the generated Siddhi apps.
@@
We can find the definition of the source stream that connects to events emitted from the IoT system via MQTT in lines 3–4 (retrieved from the YAML configuration file). Note that with our setup for the smart factory (cf. Section 2.1), we assume that each production station emits messages on one dedicated source stream. Each message contains the status of all sensors for one station at one point in time as attributes. We translate the provided patterns into the corresponding patterns of StreamingSQL referring to the source event stream(s) specified in the Radiant application. In case an activity comprises multiple stations (e.g., in the healthcare setup), thus multiple source event streams, the events from these two different streams first have to be joined and added to a new event stream to access all relevant event attributes [58]. An exemplary sink stream where detected activities events are emitted to is defined in lines 6–7. For simplicity, we only log the event occurrence, more sophisticated event sinks (e.g., using MQTT-XES [63]) can easily be added via the configuration file.
@@
Lines 9–19 in Listing LABEL:lst:siddhi-fragments contain the translated StreamingSQL queries to detect the start pattern and end pattern in the sensor data (skipping intermediate patterns) for the Sort product activity specified in Listing LABEL:lst:syntax-example2. For the start and end patterns, the changes in the attributes of two subsequent events $e1$ and $e2$ on the same event stream are analyzed for the specified conditions. We support the translation of all conditions shown in Table 1 into the concrete StreamingSQL syntax. If all conditions specified in one query are met, a new high-level event is inserted into a new event stream (here for detected patterns). Note that in line 22 we can find a more complex, automatically generated query which tracks the occurrence of all specified patterns in the sequence from start to end. With this we track the progress of the pattern detection and also ensure that the event indicating the end of an activity execution is only emitted when corresponding start and all the previous intermediate patterns have occurred in their specified sequence. Emitting the end event for an activity execution solely based on the end pattern defined in Radiant would potentially lead to incorrect detections and ambiguities as the specific pattern in the sensor events might occur more often [5, 57]. The event selected in line 23 contains all the information that we ultimately aim to derive.
@@ 4.6 Architecture of Runtime System
Figure 6 shows the overall architecture of the CEP-based process activity detection system at runtime. The Radiant applications are translated into Siddhi apps, which are deployed and executed on the service-based CEP platform Siddhi serving as light-weight runtime environment. In our proposed architecture, one Siddhi app is capable of detecting one type of activity, emitting process-level start and end events when an activity execution has been detected based on the sensor data patterns defined in the Siddhi apps. This way, we can flexibly add new Siddhi apps for new types of activities, and easily activate/deactivate these apps for specific types of activities at runtime using the REST API of the Siddhi Runner service. Furthermore, we can leverage distributed network architectures by deploying specific activity detection apps on different light-weight CEP platforms running closer to the edge (e.g., as part of one a station of the IoT system) for local sensor (pre-)processing and only emitting abstracted, higher level events [33, 28].
@@
The IoT devices produce sensor events to be sent on one or more topics to the message broker. The Siddhi apps are subscribed to these topics on the message broker to consume and process the incoming events independently from each other according to the stream configurations generated from the external configuration file. The process-level events produced and emitted from the Siddhi apps can then be either persisted in a process event log (e.g., via a subscribed persistance or logging service) for offline process mining and/or consumed and processed in streaming process mining cases by other process mining tools and services [12]. Furthermore, we setup interactive dashboards that are subscribed to process and sensor events for visualization and exploration. Figure 7 shows a screenshot of a Grafana-based dashboard which visualizes the number of detected subsequent low-level patterns for a process activity as a bar chart (top) and the associated low-level IoT sensor data as a time series (bottom).
@@
Figure 7: Visualization of activity detections and sensors in interactive dashboard (top: number of detected patterns for an activity; bottom: corresponding IoT sensor data stream) .
@@
As the architecture is event-driven and service-based [27], other more complex IoT sensors that may need additional pre-processing steps can be easily integrated, which makes the approach also applicable to more realistic IoT settings outside lab environments. Here, a basic requirement is that the corresponding software components (services) perform a processing of the complex raw data to more abstract event and (virtual) sensor data [65] that can be considered as part of the pattern and condition specification in Radiant (e.g., the abstract event machine dropped workpiece determined by a camera-based object detection service [32]). These abstracted sensor events and corresponding event streams just need to be added to the external configuration file used in Radiant (cf. Section 4.3).
@@ 4.7 Implementation
Figure 8: Composing a Radiant application in Visual Studio Code.
Radiant has been implemented using the Langium framework which uses the Language Server Protocol [66]. We decided for Langium to have a modern, flexible platform for language development that decouples the front-end from the language implementation [67]. We have exemplary integrations with Visual Studio Code (cf. Figure 8) and NeoVim supporting the user via syntax highlighting, auto-suggestions, code completion and verifications (e.g., of sensor values being in accepted ranges) based on the configuration files. The implementation of Radiant can be found on GitHub [^2] https://github.com/ics-unisg/radiant-iot-activity-dsl and as Visual Studio Code extension [^3] https://marketplace.visualstudio.com/items?itemName=mahgoh.radiant. We have implemented a code generator to translate the Radiant applications into Siddhi apps [64]. As Langium has a strong focus on extensibility, new generators to translate into code for other CEP platforms (e.g., Apache Storm) can be easily added.
@@ 5 Evaluation & Discussion
The evaluation of Radiant is based on the example processes from smart manufacturing (cf. Sect. 2.1) and smart healthcare (cf. Sect. 2.2). We asked domain experts familiar with the laboratory setups to write the Radiant applications for all relevant activities in these processes. The configuration files for the two setups were provided by the responsible IoT engineers. The Radiant applications were translated into Siddhi apps and deployed to the runtime system. We executed several instances of the processes, recorded the sensor data, and logged the process-level events of the activity detections. The processes in the smart factory were orchestrated by a BPM system [68], which generates an event log of activity executions that is used as ground truth to compare the Radiant-based activity detection with. For the executions in the smart healthcare setup, we relied on a manual monitoring and logging of the activity executions to create the ground truth. The dataset with all Radiant applications, configuration files, generated Siddhi apps, IoT sensor data, and ground truth logs can be found in [69].
@@
The comparison of the activity detection logs with the ground truth was performed using the AquDem tool presented in [70]. We use comparison metrics from several categories that are supported by AquDem as shown in Tables 2 and 3. The first two, devised by Ward et al. [71], are based on the classification of detections: 1) Two Set metrics are frame-based, which means they compare individual frames from the ground truth with those from the detection and classify them into different categories (e.g., true positives, true negatives, deletions, fragmentations, mergings, etc); 2) Event Analysis metrics are based on the Two Set metrics, using them to categorize entire activity detections as correct, deleted, merged, etc. The classifications of frames and activities provided by these metric groups can then be used to provide a range of standard metrics, e.g., precision ($Precision=\frac{TP}{TP+FP}$), recall ($Recall=\frac{TP}{TP+FN}$), the F1 score ($F1=\frac{2*Precision*Recall}{Precision+Recall}$[72] and balanced accuracy ($BA=\frac{1}{2}(\frac{TP}{TP+FN}+\frac{TN}{TN+FP})$[73]. Notably, for entire activity classifications no concept of true negatives exists and balanced accuracy is therefore not reported for the Event Analysis metrics. Furthermore, cross-correlation measures the similarity between the detected and ground truth time series by determining the time shift at which the corresponding frames are most alike and then quantifying that similarity [74]. Finally, the normalized Damerau-Levenshtein distance measures the dissimilarity between the ground truth and detection activity sequence in number of edits that would be necessary to make the sequences equal [75], normalized by the length of the longer sequence.
5.1 Smart Manufacturing
@@
We replayed the execution of five recorded IoT sensor logs, spanning in total 181.4 minutes with 424 process activity instances executed. Relevant evaluation metrics of the detection performance can be found in Table 2. A breakdown of F1 scores for the different activities is shown in Figure 9. Note that the F1 scores in Table 2, showing the overall metrics, are lower than the average over the activity-specific metrics in Figure 9, since we use micro-averaging over the activities and some of the lower performing metrics occur more frequently. Micro-averaging over the activities means that we first sum up the frame or event classifications over all activities and then calculate metrics and rates on these sums [76]. If we consider the macro-average over the activities, which does not take the number of events or frames for each activity into account, the F1 score for the Two Set metrics is $\sim 0.57$, and for the Event Analysis metrics it is $\sim 0.75$.
5.2 Smart Healthcare
For the evaluation of the healthcare scenario, we recorded and replayed the execution of 16 IoT sensor logs, each containing one process execution. These processes have been executed by students of the medical master program at our university. In total, the logs span 65.31 minutes and include 240 relevant activity executions. The detection metrics can be seen in Table 3, and F1 scores per activity in Figure 10. Note again the effect of micro-averaging when comparing the values between Table 3 and Figure 10. The macro-average F1 score over the activities in the healthcare scenario is $\sim 0.24$ for the Two Set metrics and $\sim 0.52$ for the Event Analysis metrics.

### v2

@@ 22 In SM_1 sensor o6_valve
In Listing LABEL:lst:syntax-example1 we present a Radiant example for parts of the blood donation process (cf. Section 2.4). The start of the Sanitize hands activity (line 2) is detected when the load cell sensor in the central hygiene station (CHS) changes its state from low to high (lines 4–5). This pattern requires a discretization of the sensor values emitted from the load cell (in gramms) to more abstract states. In our setup, we assume that the bottle with the hand sanitizer is placed on the load cell and it has an initial weight which will decrease with every usage. We denote this state with slowly decreasing weights as low. When the healthcare worker applies pressure to dispense liquid from the sanitizer, we can observe a significant brief peak in the load cell’s values (>> 1000 g) in a value range we denote as high. Similarly, we detect the end of this activity based on a change from high to low (lines 6–8). Note that for both conditions, we specify a time window constraint of 30 seconds as we assume that the healthcare worker might use the dispenser several times to perform one instance of the Sanitize hands activity.
@@ 4.5 Code Generation
The core requirement of the target language and execution platform should be its ability to process sensor data from one or more event streams and detect specific patterns in the current events as well as in subsequent events. Complex event processing (CEP) is the most suitable technology here as it supports exactly these features [etzion2010event]. As a concrete CEP platform, we decided for Siddhi which is part of the WSO2 stream processing platform [suhothayan2011siddhi]. Siddhi is a light-weight Java-based service optimized for high-velocity and high-throughput event processing [perera2014solving]. It features its own event processing language called StreamingSQL [jain2008towards], which is the main target (host) language of our implemented code generator to translate the Radiant artifacts to. StreamingSQL supports all concepts required to detect patterns in sensor event streams w.r.t. the corresponding activity executions.
As pointed out in [rosa-bilbao_model-driven_2022], CEP languages are still too complex and require technical expertise to effectively write event processing applications. Here, following a model-driven engineering approach [rosa-bilbao_model-driven_2022], Radiant introduces necessary abstractions to support domain experts with writing applications for the specific use case of activity detection. The Siddhi CEP platform includes a runtime system where the generated Siddhi apps can be directly deployed to and executed via a REST API, thus allowing seamless development workflows from composition with Radiant to execution and activity detection at runtime from live IoT sensor data. If necessary, domain experts can also inspect and modify the generated Siddhi apps before deployment in a web-based editor integrated with the Siddhi CEP platform.
@@
Listing LABEL:lst:siddhi-fragments presents exemplary fragments of a Siddhi app generated from the Radiant example for the Sort product activity in the production process (cf. Listing LABEL:lst:syntax-example2). We generate one app per activity that is defined in a Radiant application with the process name providing the execution context of an activity. We decided for a one-to-one mapping of activity to Siddhi app as this leads to cohesive, standalone apps that do not have other dependencies. This granularity is driven by the problem domain–the detection of activities–and allows us to easily extend the set of supported activities. Note that this corresponds to the overarching goal of this work related to the event abstraction problem that precedes process mining [diba2020extraction]. The analysis of the interrelations of detected process events and activities is subject to process mining analysis [van2012process], which is out of scope of this work. For each station with its sensors used in a Radiant application, we additionally retrieve the IoT configuration parameters from the YAML file (cf. Section 4.3) during code generation to add these configurations to the beginning of the generated Siddhi apps.
@@
We can find the definition of the source stream that connects to events emitted from the IoT system via MQTT in lines 3–4. Note that with our setup for the smart factory (cf. Section 2.3), we assume that each production station emits messages on one dedicated source stream. Each message contains the status of all sensors for one station at one point in time as attributes. We translate the provided patterns into the corresponding patterns of StreamingSQL referring to the source event stream(s) specified in the Radiant application. In case an activity comprises multiple stations (e.g., in the healthcare setup), thus multiple source event streams, the events from these two different streams first have to be joined and added to a new event stream to access all relevant event attributes [jain2008towards]. An exemplary sink stream where detected activities events are emitted to is defined in lines 6–7. For simplicity, we only log the event occurrence, more sophisticated event sinks (e.g., using MQTT-XES [burattin2020mqtt]) can be added via the configuration file.
@@
Lines 9–19 in Listing LABEL:lst:siddhi-fragments contain the translated StreamingSQL queries to detect the start pattern and end pattern in the sensor data (skipping intermediate patterns) for the Sort product activity specified in Listing LABEL:lst:syntax-example2. For the start and end patterns, the changes in the attributes of two subsequent events $e1$ and $e2$ on the same event stream are analyzed for the specified conditions. We support the translation of all conditions shown in Table 1 into the concrete StreamingSQL syntax. If all conditions specified in one query are met, a new high-level event is inserted into a new event stream (here for detected patterns). Note that in line 22 we can find a more complex, automatically generated query which tracks the occurrence of all specified patterns in the sequence from start to end. With this we track the progress of the pattern detection and also ensure that the event indicating the end of an activity execution is only emitted when corresponding start and all the previous intermediate patterns have occurred in their specified sequence. The event selected in line 23 contains all the information that we ultimately aim to derive.
@@ 4.6 Architecture of Runtime System
Addressing the research question RQ2, we present a software architecture that enables the event abstraction from IoT data at runtime. Figure 7 shows the overall architecture of the CEP-based process activity detection system at runtime. The Radiant applications are translated into Siddhi apps, which are deployed and executed on the service-based CEP platform Siddhi serving as light-weight runtime environment. In our proposed architecture, one Siddhi app is capable of detecting one type of activity, emitting process-level start and end events when an activity execution has been detected. This way, we can flexibly add new Siddhi apps for new types of activities, and easily activate/deactivate these apps for specific types of activities at runtime using the REST API of the Siddhi Runner service. Furthermore, we can leverage distributed network architectures by deploying specific activity detection apps on different light-weight CEP platforms running closer to the edge (e.g., as part of one a station of the IoT system) for local sensor pre-processing and only emitting abstracted, higher level events [seiger5165943online, systa2023liquidai]. As the main goal of our work is to provide an abstraction on top of IoT data in the form of activity executions, we expect that the number of process activities to be detected simultaneously does not need to scale significantly and that there is no need for a more fine-grained partitioning.
@@
The IoT devices produce sensor events to be sent on one or more topics to the message broker. The Siddhi apps are subscribed to these topics to consume and process the incoming events independently from each other according to the stream configurations generated from the external configuration file. The process-level events produced and emitted from the Siddhi apps can then be either persisted in a process event log (e.g., via a subscribed persistence or logging service) for offline process mining and/or consumed and processed in streaming process mining cases by other process mining tools and services [burattin2022streaming]. In case of running the event processing on an edge device with sensors directly attached, the message broker might be replaced by a more light-weight system/way to ingest the sensor data locally. Furthermore, we setup interactive dashboards that are subscribed to process and sensor events for visualization and exploration. Figure 8 shows an annotated screenshot of a Grafana-based dashboard which visualizes the number of detected subsequent low-level patterns for a process activity as a bar chart (top) and the associated low-level IoT sensor data as a time series (bottom). In the example, changes in the distance sensors followed by changes in the weight values of the load cell sensor triggered multiple patterns that are part of the Sanitize hands activity.
@@
Figure 8: Visualization of activity detections and sensors in interactive dashboard, with annotations for activity detections (top) and corresponding IoT data streams (bottom).
@@
As the architecture is event-driven and service-based [seiger2023data], other more complex IoT sensors that may need additional pre-processing steps can be integrated, which makes the approach also applicable to more realistic IoT settings outside lab environments. Here, a basic requirement is that the corresponding software components (services) perform a processing of the complex raw data to more abstract event and (virtual) sensor data [martin2021virtual] that can be considered as part of the pattern and condition specification in Radiant (e.g., the abstract event machine dropped workpiece determined by a camera-based object detection service [malburg2021object]). These abstracted sensor events and corresponding event streams just need to be added to the external configuration file (cf. Section 4.3).
@@ 4.7 Implementation
Radiant has been implemented using the Langium framework which uses the Language Server Protocol [bork2023language]. We decided for Langium to have a modern, flexible platform for language development that decouples the front-end from the language implementation [bunder2019decoupling]. We have exemplary integrations with Visual Studio Code and NeoVim supporting the user via syntax highlighting, auto-suggestions, code completion and verifications based on the configuration files. The implementation of Radiant can be found on GitHub [^4] https://github.com/ics-unisg/radiant-iot-activity-dsl and as Visual Studio Code extension [^5] https://marketplace.visualstudio.com/items?itemName=mahgoh.radiant. We have implemented a code generator to translate the Radiant applications into Siddhi apps [suhothayan2011siddhi]. As Langium has a strong focus on extensibility, new generators to translate into code for other CEP platforms (e.g., Apache Storm) can be added.
@@ 5 Evaluation & Discussion
The evaluation of the Radiant applications and the CEP-based activity detection using these applications is based on the example processes from smart manufacturing (cf. Sect. 2.3) and smart healthcare (cf. Sect. 2.4). The goal of this evaluation is to demonstrate the activity detections based on CEP apps generated from Radiant applications that were composed by domain experts for the two IoT scenarios. Through this, we aim to demonstrate the feasibility and validity of Radiant and the corresponding architecture to be used as means for IoT-based process event abstraction at runtime.
5.1 Experimental Setup
@@
To demonstrate the validity and feasibility of the DSL and its runtime capabilities, we asked domain experts familiar with the laboratory setups to write the Radiant applications for all relevant activities in the scenario processes (cf. Section 2). Both the expert from the medical domain and the expert from the manufacturing domain have been involved in the discussions of important concepts and patterns to be integrated into Radiant in earlier development stages (cf. Section 2.2). They got a brief introduction to the final language concepts and were then tasked with specifying the activity-related patterns. Both experts did not have prior skills in programming or other forms of querying/pattern languages, but they did have a basic understanding of computer science concepts, which was helpful, e.g., with correctly specifying conjunctions and disjunctions of patterns. The YAML-based configuration files for the two setups were provided by the responsible IoT engineers who are familiar with the laboratory setups and their existing sensors. The Radiant applications were translated into Siddhi apps and deployed to the runtime system. We executed several instances of the processes, recorded the sensor data, and logged the process-level events of the activity detections. The processes in the smart factory were orchestrated by a BPM system [seiger2022integrating], which generates an event log of activity executions that is used as ground truth to compare the Radiant-based activity detection with. For the executions in the smart healthcare setup, we relied on a manual monitoring and logging of the activity executions to create the ground truth. The dataset with all Radiant applications, configuration files, generated Siddhi apps, IoT sensor data, and ground truth logs can be found in [seiger_2025_15773369].
5.2 Performance
The Siddhi CEP engine has proven its capabilities of filtering and processing more than one million events per second in memory and more than 300.000 events per second via network [perera2014solving]. These benchmarks make the engine a suitable technology for larger scale, distributed IoT deployments and event abstraction at runtime.
During our experiments capturing, replaying and processing the IoT data streams, we did not experience any performance-related issues on state-of-the-art laptop computers. We were able to run 14 different CEP applications for activity detection in parallel in our laboratory setup, processing 1200 IoT sensor readings per minute (limited by the sampling frequency of the sensors) [seiger5165943online]. This confirms that our current choices of granularity (i.e., one Siddhi app per activity) and one instance per app on the CEP platform are feasible in the given IoT setups.
An exemplary runtime resource consumption on our experimentation laptop (Intel Core i7-8550U@1.8GHZ, 4 Cores, 16 GB RAM, Windows 11) running the Java-based Siddhi Runner (v5.1.2666https://siddhi.io/en/v5.1/download/) application in a 64-bit JVM (Java 1.8) with 14 active Siddhi apps (generated from Radiant) included CPU loads between 0.2% and 1% and main memory consumption of 9.4 MB per Siddhi app with active connections to 31 sensors via MQTT; each sensor sampling its data at 2 Hz. The memory consumption of the Siddhi Runner including the JVM it runs in, without any active CEP apps was 224 MB on average. Latency measurements resulted in average event processing times of less than 50 ms, which include sampling a sensor, sending its data to an MQTT broker on a different computer via Gigabit-Ethernet, consuming it from the MQTT broker via the corresponding Siddhi app, processing the sensor data, and logging the occurrence of a process-level event. Large-scale performance benchmarks with high-frequency IoT sensors and devices remain subject to future work.
5.3 Detection Metrics
@@
The comparison of the activity detection logs with the ground truth was performed using the AquDem tool presented in [kurz2024activity]. We use comparison metrics from several categories that are supported by AquDem as shown in Tables 2 and 3. The first two, devised by Ward et al. [wardmetr2011], are based on the classification of detections: 1) Two Set metrics are frame-based, which means they compare individual frames from the ground truth with those from the detection and classify them into different categories (e.g., true positives, true negatives, deletions, fragmentations, mergings, etc); 2) Event Analysis metrics are based on the Two Set metrics, using them to categorize entire activity detections as correct (i.e., a detected event corresponds to the ground truth event), deleted (i.e., an event from the ground truth was not detected), fragmented (i.e., one event in the ground truth is recognized by several detected events), merged (i.e., several ground truth events correspond to one detected event), etc [wardmetr2011]. The classifications of frames and activities provided by these metric groups can then be used to provide a range of standard metrics, e.g., precision ($Precision=\frac{TP}{TP+FP}$), recall ($Recall=\frac{TP}{TP+FN}$), the F1 score ($F1=\frac{2*Precision*Recall}{Precision+Recall}$[Taha2015] and balanced accuracy ($BA=\frac{1}{2}(\frac{TP}{TP+FN}+\frac{TN}{TN+FP})$[mosley2013balanced]. Notably, for entire activity classifications no concept of true negatives exists and balanced accuracy is therefore not reported for the Event Analysis metrics. Furthermore, cross-correlation measures the similarity between the detected and ground truth time series by determining the time shift at which the corresponding frames are most alike and then quantifying that similarity [lyon2010discrete]. Finally, the normalized Damerau-Levenshtein distance measures the dissimilarity between the ground truth and detection activity sequence in number of edits that would be necessary to make the sequences equal [damerau_technique_1964], normalized by the length of the longer sequence.
5.4 Experiments and Results: Smart Manufacturing
@@
We replayed the execution of five recorded IoT sensor logs, spanning in total 181.4 minutes with 424 process activity instances executed. These logs contain the timestamped readings from 77 distinct sensors and actuators, each sampled at 0.5 Hz. Activities in the smart factory include the transportation of workpieces, storage and unloading, sorting, simulated burning and milling (cf. Figure 9). Their durations were in the range of 5 seconds (Calibrate VGR) to 49 seconds (Unload from High-bay Warehouse). Relevant evaluation metrics of the detection performance can be found in Table 2. A breakdown of F1 scores for the different activities is shown in Figure 9. Note that the F1 scores in Table 2, showing the overall metrics, are lower than the average over the activity-specific metrics in Figure 9, since we use micro-averaging over the activities and some of the lower performing metrics occur more frequently. Micro-averaging over the activities means that we first sum up the frame or event classifications over all activities and then calculate metrics and rates on these sums [sokolova_systematic_2009]. If we consider the macro-average over the activities, which does not take the number of events or frames for each activity into account, the F1 score for the Two Set metrics is $\sim 0.57$, and for the Event Analysis metrics it is $\sim 0.75$.
5.5 Experiments and Results: Smart Healthcare
For the evaluation of the healthcare scenario, we recorded and replayed the execution of 16 IoT sensor logs, each containing one process execution. These processes have been executed by students of the medical master program at our university. In total, the logs span 65.31 minutes and include 240 relevant activity executions. The IoT logs contain 31 distinct, timestamped sensor values, each sampled at 2 Hz. Activities in the smart healthcare setup include the check-ins of patient and healthcare worker, hand hygiene, applying/remove the tourniquet, inserting/removing the needle and operating the blood drawing machine (cf. Figure 10). Their durations were in the range of 3 seconds (Donor check-in) to 20 seconds (Sanitize hygiene). The detection metrics can be seen in Table 3, and F1 scores per activity in Figure 10. Note again the effect of micro-averaging when comparing the values between Table 3 and Figure 10. The macro-average F1 score over the activities in the healthcare scenario is $\sim 0.24$ for the Two Set metrics and $\sim 0.52$ for the Event Analysis metrics.